Business and finance. The cost in dollars of manufacturing w wing nuts is given by the expression 0.07w + 13.3. Find the cost when 375 wing nuts are made. What is the average cost to manufacture one wing nut?
Answer by checkley75(3666)

0.07W+13.3
0.07*375+13.3=26.25+13.3=39.55 COST OF 375 WING NUTS.
39.55/375=.105466 COST OF EACH WING NUT.